OpenStudy (anonymous):

is my answer correct just reply yes or no thx. A circle has a radius of 12 centimeters. If 5/6 of the circle is cut away, what is the area of the remaining portion? Use 3.14 75.36 cm^2 or 75.4 whichever

OpenStudy (anonymous):

just tell me this it say 5/6 i calculated each angle would be 60 that's how i went about to figure it out each angle wouldn't be this?

OpenStudy (saifoo.khan):

Hey, sorry. you are correct.

OpenStudy (anonymous):

im correct about each angle being 60 or my answer is accurate?

OpenStudy (saifoo.khan):

Answer. 75.4

OpenStudy (anonymous):

ok thx

OpenStudy (saifoo.khan):

\[\frac{60}{360 } \times \pi (12)^2\]